OpenStudy (anonymous):

Rewrite in simplest radical form x^5/6 divided by x^1/6

Nnesha (nnesha):

when we divide same base , `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (anonymous):

Do you subtract both the bottom number and the top number or just the top one?

Nnesha (nnesha):

just subtract their exponents top exponent `minus` bottom exponents base would stay the same

OpenStudy (anonymous):

.67

Nnesha (nnesha):

hmm we need to write it in radical form so keep the fraction

OpenStudy (anonymous):

How?

Nnesha (nnesha):

how did you get .67 ??

Nnesha (nnesha):

i mean what did you subtract to get .67 what did you type into the calculator ?

OpenStudy (anonymous):

I subtract 5/6 from 1/6

Nnesha (nnesha):

yes so \[\huge\rm \frac{ 5 }{ 6 }-\frac{ 1 }{ 6 }\] what's the common denominator ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

x^4/6?

Nnesha (nnesha):

yes right so since the denominator of both fraction are the same \[\frac{ 5-1 }{ 6 }\] so you can just subtract the numbers in the numerator

Nnesha (nnesha):

yes right reduce the fraction

OpenStudy (anonymous):

x^2/3?

Nnesha (nnesha):

yes right now apply the exponent rule to convert into radical form \[\huge\rm \sqrt[n]{x^m}=x^\frac{ m }{ n}\]

OpenStudy (anonymous):

\[\sqrt[3]{x ^{2}}\]

Nnesha (nnesha):

looks good